OpenStudy (anonymous):

Check my answer please ! What is the slope of a line parallel to the line whose equation is 3x – 4y = 8? (Points : 4) (Wont let me add the fractions ) i got 4/3

OpenStudy (ybarrap):

Can you put your equation in the following form? y=mx+b If you can, your answer is the "m" in this equation.

OpenStudy (anonymous):

How do I get it in that form?

OpenStudy (anonymous):

no,it is not 4/3

OpenStudy (anonymous):

its 3/4

OpenStudy (anonymous):

correct.

OpenStudy (ybarrap):

$$ 3x – 4y = 8\\ 3x – 4y\color{red}{-3x}=8\color{red}{-3x}\\ -4y=8-3x\\ $$ Do you see this? Now divide the whole thing by -4 and see what you get for the slope $$ \cfrac{-4y}{-4}=\cfrac{8-3x}{-4} $$ See it now? y = mx + b